OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (7, -3) and (4, -8)

OpenStudy (sleepyjess):

let's put it in point slope form first

OpenStudy (anonymous):

How do you do that?

OpenStudy (sleepyjess):

do you know what point slope form is?

OpenStudy (anonymous):

No, I haven't really worked with point slope

OpenStudy (sleepyjess):

Wait, no, maybe not point slope. Have you learned slope intercept form?

OpenStudy (anonymous):

Yes ma'am. y=mx+b

OpenStudy (sleepyjess):

Okay, let's use that then :) First we need to find the slope

OpenStudy (anonymous):

Alright, gimme a second.

OpenStudy (sleepyjess):

ok :)

OpenStudy (anonymous):

5/3

OpenStudy (sleepyjess):

Yes, now we can use that to find the y intercept

OpenStudy (anonymous):

-44/3

OpenStudy (sleepyjess):

what is -44/3?

OpenStudy (anonymous):

The y-intercept.

OpenStudy (sleepyjess):

the y-intercept is where x = 0

OpenStudy (anonymous):

Oh..lol I just went to Mathway.com and got the slope and y-intercept from there

OpenStudy (sleepyjess):

do you know how to find slope?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (sleepyjess):

Ok, I see what you are saying, the y intercept is 0, 44/3

OpenStudy (sleepyjess):

Now we have y = 5/3x - 44/3

OpenStudy (sleepyjess):

I am just confusing myself now...

OpenStudy (anonymous):

Me too! Lol

OpenStudy (sleepyjess):

@hartnn

hartnn (hartnn):

whats the standard form of line ?

OpenStudy (anonymous):

I don't know! That's what I'm trying to figure out lol

hartnn (hartnn):

its ax+by =c

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

y = 5/3x - 44/3 multiply throughout by 3

OpenStudy (anonymous):

Even y?

hartnn (hartnn):

everything that will eliminate fractions on right side and make our life easier ^_^

OpenStudy (anonymous):

3y=5x-44

OpenStudy (anonymous):

Would the answer be 5x+3y=-44?

OpenStudy (sleepyjess):

Almost

OpenStudy (sleepyjess):

You added 5x instead of subtracting

OpenStudy (anonymous):

5x was never negative though

OpenStudy (sleepyjess):

but we need to subtract to move it to the left side

OpenStudy (anonymous):

-5x+3y=-44 ??

OpenStudy (sleepyjess):

yes, and personally, I would probably multiply by -1 to make x positive, but it isn't required

hartnn (hartnn):

lets do that, it wil be more accurate

OpenStudy (anonymous):

???

hartnn (hartnn):

keeping the co-efficient of x as positive

hartnn (hartnn):

multiply by -1 throughout

OpenStudy (anonymous):

So it would make 3y negative

OpenStudy (sleepyjess):

Everything just changes signs

OpenStudy (anonymous):

So the answer is 5x-3y=44

OpenStudy (sleepyjess):

Yes

OpenStudy (anonymous):

AHH THANK YOU BOTH

OpenStudy (anonymous):

This is all so confusing to me. I have a few more questions if y'all wouldn't mind helping me

OpenStudy (sleepyjess):

you're welcome sure :)